OpenStudy (anonymous):

find the exact value of the expression. help please! sin[-11pi/12]

OpenStudy (anonymous):

You want to convert to degrees?

OpenStudy (anonymous):

no I believe we have to break the -11pi/12 to two angles that we can find their value and go from there. but I am not entirely sure as that's only what the instructions say.

OpenStudy (mertsj):

\[\sin \frac{x}{2}=\sqrt{\frac{1-\cos x}{2}}\]

OpenStudy (mertsj):

Use 1=-11 pi/6 for x then x/2 is -11 pi/12

OpenStudy (anonymous):

 sin (-165)=sin(-15) =-sin(15) =-sin(60-45) and go from there